OpenStudy (anonymous):

5/6=-10/3y-5/9 what does y equal ?? thanks!!

OpenStudy (akshay_budhkar):

-12/5....

OpenStudy (anonymous):

10/3y - 5/9 = 5/6 10 - 5 = 5 3y - 9 = 6 focusing on 3y - 9 = 6 3y = 15 divide both sides by 3 y = 5

OpenStudy (mimi_x3):

5/6 = -(30-5y)/9Y 45y = 6(-30-5y) 45y = 6(-30-5y) 75y = -180 - 30y 75y = -180 y = -12/5 (not sure if im right though)

OpenStudy (akshay_budhkar):

its right mimi

OpenStudy (mimi_x3):

yay! i got it right xDD

OpenStudy (akshay_budhkar):

zhill you made a mistake in copying the sum